Grade 11IIT JEE Entrance Exam

two tangents ae drawn from the point (1,4)to the parabola y^2=4ax.Angle between tangents is :

aditya malviya
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
 
Angle is 60 degrees
As y2=4x, so a=1
the equation of tangent is y=mx+1/m.....(i)
it passes through (1,4).
so the equation (i) becomes m2-4m+1=0
solving the quadratic and substituting its value in formula tan theta=m2-m/1+m2m, you will get your answer.